OpenStudy (anonymous):

FAN MEDAL AND TESTIMONY ONLY ONE QUESTION!!!! Two equations are given below: a − 3b = 4 a = b − 2 What is the solution to the set of equations in the form (a, b)? (−2, −2) (−3, −1) (−9, −7) (−5, −3)

Vocaloid (vocaloid):

a = 3b + 4 a = b - 2 3b + 4 = b - 2 solve for b